# Using Dynamic Safety Margins as Control Barrier Functions

### v4

@@ 3.1 DSMs are CBFs
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$ and let $\alpha\in\mathscr{K}$ be such that (9) holds. If $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ is such that
$\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}=0,$ (30)
then, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
Proof 3.14.
@@
Note that since $\big(\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, it follows that $\forall i\in\{1,\ldots,n_{c}\}$,
$\dot{\Delta}_{i}\big(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big)\!$ $=\!\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\!+\!\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})\!&\!\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}\!\begin{bmatrix}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ 0\end{bmatrix}$
$=\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\geq-\alpha\big(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big).$
@@
Therefore, constraint (28) is trivially satisfied at $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for any augmented input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})$. We conclude that $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
3.2 DSM-CBF Optimization-based Policy
@@
Recalling Theorem 3.9, it is possible to reformulate the CBF policy $\beta$ for the augmented system $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. To do so, let $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ be a desired reference, let $\kappa:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$ be a nominal control policy with desirable closed-loop performance, and let $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ be a navigation field. The following theorem formulates the CBF-based policy $\beta$ for DSMs and the augmented system (24).
Theorem 3.15 (DSM-CBF policy).
@@
Let $\eta>0$ be a constant scalar. If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then there exists $\alpha\in\mathscr{K}$ such that the optimization-based policy $\beta:\tilde{\mathcal{C}}\to\mathcal{U}\times\mathbb{R}^{l}$
@@
$\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\!\underset{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})}{\operatorname{argmin}}\|\boldsymbol{\mathbf{u}}\!-\!\kappa(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\|^{2}\!+\eta\|\boldsymbol{\mathbf{w}}\!-\!\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\|^{2},$ (31)
@@
is feasible for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$. Moreover, if $\beta$ is locally Lipschitz continuous and 0 is a regular value of each $\Delta_{i}$, then the trajectory of the augmented system (24) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big(\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big)\in\tilde{\mathcal{C}}$.
Proof 3.16.
@@
Feasibility of the DSM-CBF policy follows by Corollary 3.11. Any locally Lipschitz controller in $\tilde{\mathcal{K}}_{\alpha}$ certifies the local (in time) existence and uniqueness of the solution to the closed-loop system [37]. Due to the the invariance of $\tilde{\mathcal{C}}$, there can be no finite-time escape instances and the existence and uniqueness holds for all $t\geq 0$. Safety follows from $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$.
Remark 3.17.
@@
Convexity of the set $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, and therefore problem (31), is linked to convexity of $\mathcal{U}$. In particular, if the input constraint set $\mathcal{U}$ is polyhedral, the optimization problem (31) becomes a quadratic program (QP) for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, which can be solved efficiently.
Remark 3.18.
@@
Remark 3.19.
If $\mathcal{V}$ is convex, a suitable navigation field is $\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})=\boldsymbol{\mathbf{r}}-\boldsymbol{\mathbf{v}}$. Otherwise, please refer to works like [38].
Remark 3.20.
Note that the nominal controller $\kappa$ and the prestabilizing controller $\pi$ can be different. In practice, we observed that mild prestabilizing controllers $\pi$ lead to larger feasible regions on the CBF-based program, allowing the safe input to be closer to the nominal target $\kappa$.
3.3 Existence and Uniqueness of Solutions
@@
Given any initial condition $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, the optimization-based policy $\beta$ for DSM-based CBFs, defined in (31), is always feasible. If, additionally, it is locally Lipschitz continuous, then the trajectory for the closed-loop system
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})\\ 0\end{bmatrix}+\begin{bmatrix}g(\boldsymbol{\mathbf{x}})&0\\ 0&I\end{bmatrix}\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (32)
@@
Theorem 3.21 ([40, Theorem 5.2]).
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$, let $\rho$ be a navigation field and let $\eta>0$. Assume the input constraints are polyhedral $\mathcal{U}=\{\boldsymbol{\mathbf{u}}\in\mathbb{R}^{m}:M\boldsymbol{\mathbf{u}}\leq\boldsymbol{\mathbf{b}}\}$, and the functions $\pi(\cdot,\boldsymbol{\mathbf{r}})$ and $\rho(\cdot,\boldsymbol{\mathbf{r}})$ are continuously differentiable. If these constraint qualifications (CQs) hold
@@
    (Slater CQ) There exists $\alpha\in\mathscr{K}$ such that for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\operatorname{Int}\left(\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\right)\neq\emptyset$;
@@
    (Constant Rank CQ) There exists a neighborhood of every pair $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ wherein, for any subset $\mathcal{I}\subset\mathcal{J}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, where $\mathcal{J}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is the set of active constraints, the number of linearly independent vectors in $\left\{\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix},i\in\mathcal{I}\right\}$ remains constant;
@@
Remark 3.22.
In the special case that $n_{c}=1$, both Slater and Constant Rank CQs hold if $\Delta$ has uniform relative degree $1$ over $\tilde{\mathcal{C}}$ and $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\operatorname{Int}\left(\mathcal{U}\right)$ for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$.
4 Lyapunov-based Dynamic Safety Margins
@@
Definition 4.23.
A continuously differentiable function $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ is a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ if there exist functions $\gamma_{1},\gamma_{2}\in\mathscr{K}$ and $\forall\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, there exists a neighborhood of $\bar{x}(\boldsymbol{\mathbf{v}})$, denoted $\mathcal{D}_{\boldsymbol{\mathbf{v}}}\subset\mathbb{R}^{n}$, wherein
$\gamma_{1}\big(\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big)$ $\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq\gamma_{2}\big(\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|\big),$ (33a)
$\frac{\partial V}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ $\leq 0.$ (33b)
Remark 4.24.
Note that the reference-dependent Lyapunov function $V$ needs not certify asymptotic stability of the equilibria of the prestabilized system $f_{\pi}$. That is, $\dot{V}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$ is acceptable for $\boldsymbol{\mathbf{x}}\neq\bar{x}(\boldsymbol{\mathbf{v}})$. This allows for more general classes of Lyapunov functions to be considered and ultimately simplifies the design of the DSM.
@@
Given any $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$, $\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\setminus\{\bar{x}(\boldsymbol{\mathbf{v}})\}$ is a regular point of $V(\cdot,\boldsymbol{\mathbf{v}}):\mathbb{R}^{n}\to\mathbb{R}$.
Recalling the reference-dependent state constraint set $\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ defined in (19), let $\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}=\mathbb{R}^{n}\setminus\mathcal{X}_{\boldsymbol{\mathbf{v}}}$ be its complement, which is open in $\mathbb{R}^{n}$. We define the safety threshold value $\Gamma^{*}:\mathbb{R}^{l}\to\mathbb{R}$
$\Gamma^{*}(\boldsymbol{\mathbf{v}})\triangleq\inf_{\boldsymbol{\mathbf{x}}\in\>\mathcal{X}^{c}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (34)
When positive, $\Gamma^{*}(\boldsymbol{\mathbf{v}})$ represents the largest constraint-admissible level set of the Lyapunov function $V$ at reference $\boldsymbol{\mathbf{v}}$. Our definition for $\Gamma^{*}$ is an extension of the threshold value in [24] to the entire reference space $\mathbb{R}^{l}$. With this, $\Gamma^{*}$ has the following property.
Lemma 4.25.
Let $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ be a reference-dependent Lyapunov function that satisfies Assumption 4. We have that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0$ if and only if $\boldsymbol{\mathbf{v}}\in\mathcal{V}$.
Proof 4.26.
Let $\boldsymbol{\mathbf{v}}\in\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Recalling that $\forall\boldsymbol{\mathbf{x}}\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, $V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, we conclude that $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq 0$.
For the other direction, we show the contrapositive statement. Let $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}\setminus\mathcal{V}$ be given. Then, $\bar{x}(\boldsymbol{\mathbf{v}})\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$, which implies that $\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$. Since $\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$ is open in $\mathbb{R}^{n}$, there exists $\delta>0$ such that $\mathcal{B}\big(\bar{x}(\boldsymbol{\mathbf{v}}),\delta\big)=\{\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}\mid\|\boldsymbol{\mathbf{x}}-\bar{x}(\boldsymbol{\mathbf{v}})\|<\delta\}\subset\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}$. Thus,
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=-\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq-\gamma_{1}(\delta)<0.$ (35)
@@
$\overline{\Gamma}(\boldsymbol{\mathbf{v}})\triangleq\inf_{\boldsymbol{\mathbf{x}}\in\partial\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}).$ (36)
@@
Theorem 4.27 (Lyapunov-based DSMs).
@@
$\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}\Gamma^{*}(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ (1-\epsilon)\overline{\Gamma}(\boldsymbol{\mathbf{v}})-V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix}$, (37)
is a dynamic safety margin for $\pi$.
Proof 4.28.
@@
To prove property (22a), let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{D}}$. Then, $\Gamma^{*}(\boldsymbol{\mathbf{v}})\geq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$, which implies that $\boldsymbol{\mathbf{v}}\in\mathcal{V}$ by Lemma 4.25. Next, for a contradiction, assume $\boldsymbol{\mathbf{x}}\notin\mathcal{X}_{\boldsymbol{\mathbf{v}}}$. Then
$\Gamma^{*}(\boldsymbol{\mathbf{v}})=\inf_{\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}^{c}\cap\mathcal{D}_{\boldsymbol{\mathbf{v}}}}V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq V(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$

### v5

@@ III-A DSMs are CBFs
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM and let $\alpha\in\mathscr{K}$ be such that (25) holds. If $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$ is such that
$\begin{bmatrix}\frac{\partial\Delta}{\partial\bm{x}}g(\bm{x})&\frac{\partial\Delta}{\partial\bm{v}}\end{bmatrix}=0,$ (27)
then, $\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})=\mathcal{U}\times\mathbb{R}^{l}$.
Proof:
@@
Note that since $\big(\pi(\bm{x},\bm{v}),0\big)\in\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})$, it follows that $\forall i\in\{1,\ldots,n_{c}\}$,
$\dot{\Delta}_{i}\big(\bm{x},\bm{v},\pi(\bm{x},\bm{v}),0\big)\!$ $=\!\frac{\partial\Delta_{i}}{\partial\bm{x}}f(\bm{x})\!+\!\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\bm{x}}g(\bm{x})\!&\!\frac{\partial\Delta_{i}}{\partial\bm{v}}\end{bmatrix}\!\begin{bmatrix}\pi(\bm{x},\bm{v})\\ 0\end{bmatrix}$
$=\frac{\partial\Delta_{i}}{\partial\bm{x}}f(\bm{x})\geq-\alpha\big(\Delta_{i}(\bm{x},\bm{v})\big).$
@@
Therefore, constraint (25) is trivially satisfied at $(\bm{x},\bm{v})$ for any augmented input $(\bm{u},\bm{w})$. We conclude that $\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})=\mathcal{U}\times\mathbb{R}^{l}$. ∎
III-B DSM-CBF Optimization-based Policy
@@
Recalling Theorem 4, it is possible to reformulate the CBF policy $\beta$ for the augmented system $(\bm{x},\bm{v})$. To do so, let $\bm{r}\in\mathbb{R}^{l}$ be a desired reference, let $\kappa:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$ be a nominal control policy with desirable closed-loop performance, and let $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ be a navigation field. The following theorem formulates the DSM-CBF-based policy $\tilde{\beta}$ for the augmented system (21).
Theorem 5 (DSM-CBF policy)
@@
Let $\eta>0$ be a constant scalar. If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then there exists $\alpha\in\mathscr{K}$ such that the optimization-based policy $\tilde{\beta}:\tilde{\mathcal{C}}\to\mathcal{U}\times\mathbb{R}^{l}$
@@
$\tilde{\beta}(\bm{x},\bm{v})\!=\!\!\!\!\underset{(\bm{u},\bm{w})\in\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})}{\operatorname{argmin}}\|\bm{u}\!-\!\kappa(\bm{x},\bm{r})\|^{2}\!+\eta\|\bm{w}\!-\!\rho(\bm{v},\bm{r})\|^{2},$ (28)
@@
is feasible for all $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$. Moreover, if $\tilde{\beta}$ is locally Lipschitz continuous and $\Delta$ satisfies the MFCQ, then the trajectory of the augmented system (21) under policy $(\bm{u},\bm{w})=\tilde{\beta}(\bm{x},\bm{v})$ exists, is unique, and satisfies $x(t)\in\mathcal{X}$ and $v(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big(x(0),v(0)\big)\in\tilde{\mathcal{C}}$.
Proof:
@@
Feasibility of the DSM-CBF policy follows by Theorem 4. Any locally Lipschitz controller in $\tilde{\mathcal{K}}_{\alpha}$ certifies the local (in time) existence and uniqueness of the solution to the closed-loop system [28]. Recalling Corollary 1, control invariance of $\tilde{\mathcal{C}}$ implies there can be no finite-time escape instances and the existence and uniqueness holds for all $t\geq 0$. Safety follows from $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$. ∎
Remark 3
@@
Convexity of the set $\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})$, and therefore problem (28), is linked to convexity of $\mathcal{U}$. In particular, if the input constraint set $\mathcal{U}$ is polyhedral, the optimization problem (28) becomes a quadratic program (QP) for all $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$, which can be solved efficiently.
Remark 4
@@
Remark 5
If $\mathcal{V}$ is convex, a suitable navigation field is $\rho(\bm{v},\bm{r})=\bm{r}-\bm{v}$. Otherwise, please refer to works like [41].
Remark 6
Note that the nominal controller $\kappa$ and the prestabilizing controller $\pi$ can be different. In practice, we observed that mild prestabilizing controllers $\pi$ lead to larger $\tilde{\mathcal{K}}_{\alpha}$, which is the feasible region of (28), allowing the safe input to be closer to the nominal target $\kappa$.
III-C Existence and Uniqueness of Solutions
@@
As stated in Theorem 5, given any initial condition $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$, the optimization-based policy $\tilde{\beta}$ for DSM-CBFs, defined in (28), is always feasible. If, additionally, it is locally Lipschitz continuous, then the trajectory for the closed-loop system
$\begin{bmatrix}\dot{\bm{x}}\\ \dot{\bm{v}}\end{bmatrix}=\begin{bmatrix}f(\bm{x})\\ 0\end{bmatrix}+\begin{bmatrix}g(\bm{x})&0\\ 0&I\end{bmatrix}\beta(\bm{x},\bm{v}),$ (29)
@@
Theorem 6 ([32, Theorem 5.2])
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM, let $\rho$ be a navigation field and let $\eta>0$. Assume the input constraints are polyhedral $\mathcal{U}=\{\bm{u}\in\mathbb{R}^{m}\mid M\bm{u}\leq\bm{b}\}$, and the functions $\pi(\cdot,\bm{r})$ and $\rho(\cdot,\bm{r})$ are continuously differentiable. If these constraint qualifications (CQs) hold
@@
    (Slater CQ) There exists $\alpha\in\mathscr{K}$ such that for all $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$, $\operatorname{Int}\left(\tilde{\mathcal{K}}_{\alpha}(\bm{x},\bm{v})\right)\neq\emptyset$;
@@
    (Constant Rank CQ) There exists a neighborhood of every pair $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$ wherein, for any subset $\mathcal{J}\subset\mathcal{I}(\bm{x},\bm{v})$, where $\mathcal{I}(\bm{x},\bm{v})$ is the set of active constraints, the number of linearly independent vectors in $\left\{\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\bm{x}}g(\bm{x})&\frac{\partial\Delta_{i}}{\partial\bm{v}}\end{bmatrix},i\in\mathcal{J}\right\}$ remains constant;
@@
Remark 7
In the special case that $n_{c}=1$, both Slater and Constant Rank CQs hold if $\Delta$ has uniform relative degree $1$ over $\tilde{\mathcal{C}}$ and $\pi(\bm{x},\bm{v})\in\operatorname{Int}\left(\mathcal{U}\right)$ for all $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}$.
IV Lyapunov-based Dynamic Safety Margins
@@
Definition 4
A continuously differentiable function $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ is a reference-dependent Lyapunov function for the prestabilized system $f_{\pi}$ if there exist functions $\gamma_{1},\gamma_{2}\in\mathscr{K}$ and $\forall\bm{v}\in\mathbb{R}^{l}$, there exists a neighborhood of $\bar{x}(\bm{v})$, denoted $\mathcal{D}_{\bm{v}}\subset\mathbb{R}^{n}$, wherein
$\gamma_{1}\big(\|\bm{x}-\bar{x}(\bm{v})\|\big)$ $\leq V(\bm{x},\bm{v})\leq\gamma_{2}\big(\|\bm{x}-\bar{x}(\bm{v})\|\big),$ (30a)
$\frac{\partial V}{\partial\bm{x}}f_{\pi}(\bm{x},\bm{v})$ $\leq 0.$ (30b)
Remark 8
Note that the reference-dependent Lyapunov function $V$ needs not certify asymptotic stability of the equilibria of the prestabilized system $f_{\pi}$. That is, $\dot{V}(\bm{x},\bm{v})=0$ is acceptable for $\bm{x}\neq\bar{x}(\bm{v})$. This allows for more general classes of Lyapunov functions to be considered and ultimately simplifies the design of the DSM.
@@
Given any $\bm{v}\in\mathbb{R}^{l}$, $\bm{x}\in\mathcal{D}_{\bm{v}}\setminus\{\bar{x}(\bm{v})\}$ is a regular point of $V(\cdot,\bm{v}):\mathbb{R}^{n}\to\mathbb{R}$.
Recalling the reference-dependent state constraint set $\mathcal{X}_{\bm{v}}$ defined in (16), let $\mathcal{X}^{c}_{\bm{v}}=\mathbb{R}^{n}\setminus\mathcal{X}_{\bm{v}}$ be its complement, which is open in $\mathbb{R}^{n}$. We define the safety threshold value $\Gamma^{*}:\mathbb{R}^{l}\to\mathbb{R}$
$\Gamma^{*}(\bm{v})\triangleq\inf_{\bm{x}\in\>\mathcal{X}^{c}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v})-\inf_{\bm{x}\in\mathcal{X}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v}).$ (31)
When positive, $\Gamma^{*}(\bm{v})$ represents the largest constraint-admissible level set of the Lyapunov function $V$ at reference $\bm{v}$. Our definition for $\Gamma^{*}$ is an extension of the threshold value in [38] to the entire reference space $\mathbb{R}^{l}$. With this, $\Gamma^{*}$ has the following property.
Lemma 2
Let $V:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}$ be a reference-dependent Lyapunov function that satisfies Assumption 4. We have that $\Gamma^{*}(\bm{v})\geq 0$ if and only if $\bm{v}\in\mathcal{V}$.
Proof:
Let $\bm{v}\in\mathcal{V}$ be given. Then, $\bar{x}(\bm{v})\in\mathcal{X}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}$, which implies that $\inf_{\bm{x}\in\mathcal{X}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v})=0$. Recalling that $\forall\bm{x}\in\mathcal{D}_{\bm{v}}$, $V(\bm{x},\bm{v})\geq 0$, we conclude that $\Gamma^{*}(\bm{v})\geq 0$.
For the other direction, we show the contrapositive statement. Let $\bm{v}\in\mathbb{R}^{l}\setminus\mathcal{V}$ be given. Then, $\bar{x}(\bm{v})\in\mathcal{X}_{\bm{v}}^{c}\cap\mathcal{D}_{\bm{v}}$, which implies that $\inf_{\bm{x}\in\mathcal{X}_{\bm{v}}^{c}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v})=0$. Since $\mathcal{X}_{\bm{v}}^{c}\cap\mathcal{D}_{\bm{v}}$ is open in $\mathbb{R}^{n}$, there exists $\delta>0$ such that $\mathcal{B}\big(\bar{x}(\bm{v}),\delta\big)=\{\bm{x}\in\mathbb{R}^{n}\mid\|\bm{x}-\bar{x}(\bm{v})\|<\delta\}\subset\mathcal{X}_{\bm{v}}^{c}\cap\mathcal{D}_{\bm{v}}$. Thus,
$\Gamma^{*}(\bm{v})=-\inf_{\bm{x}\in\mathcal{X}_{\bm{v}}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v})\leq-\gamma_{1}(\delta)<0.$ (32)
@@
$\overline{\Gamma}(\bm{v})\triangleq\inf_{\bm{x}\in\partial\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v}).$ (33)
@@
Theorem 7 (Lyapunov-based DSMs)
@@
$\Delta(\bm{x},\bm{v})=\begin{bmatrix}\Gamma^{*}(\bm{v})-V(\bm{x},\bm{v})\\ (1-\epsilon)\overline{\Gamma}(\bm{v})-V(\bm{x},\bm{v})\end{bmatrix}$, (34)
is a dynamic safety margin.
Proof:
@@
To prove property (19a), let $(\bm{x},\bm{v})\in\tilde{\mathcal{C}}\subset\tilde{\mathcal{D}}$. Then, $\Gamma^{*}(\bm{v})\geq V(\bm{x},\bm{v})\geq 0$, which implies that $\bm{v}\in\mathcal{V}$ by Lemma 2. Next, for a contradiction, assume $\bm{x}\notin\mathcal{X}_{\bm{v}}$. Then
$\Gamma^{*}(\bm{v})=\inf_{\bm{x}\in\mathcal{X}_{\bm{v}}^{c}\cap\mathcal{D}_{\bm{v}}}V(\bm{x},\bm{v})\leq V(\bm{x},\bm{v}).$